Sixty percent of new drivers have had driver education. During their first year, new drivers without driver education have a probability of 0.08 of having an accident, but new drivers with driver education have only a 0.05 probability of an accident. What is the probability a new driver has had driver education, given that the driver has had no accident the first year?
Answer #1
Education No Education Total
Accident 0.05 * 60 = 3 0.08 * 40 = 3.2 6.2
No Accident 60 - 3 = 57 40 - 3.2 = 36.8 93.8
Total 60 40 100

P(Education|No Accident) = (57/100)/(93.8/100) = 0.6077